OpenStudy (anonymous):

4. Identify the point (x1, y1) from the equation: y – 9 = 3(x + 2) (Points : 1) (–2, 9) (9, –2) (2, –9) (–8, 2)

OpenStudy (anonymous):

y - y1 = m(x - x1) y - 9 = 3(x + 2) so x1 = -2 and y1 = 9....let me check it to make sure y - y1 = m(x - x1) slope = 3 (-2,9)...x1 = -2 and y1 = 9 sub y - 9 = 3(x - (-2) y -9 = 3(x + 2) (correct) answer is (-2,9)